### v1

# AI-Driven Cytomorphology Image Synthesis for Medical Diagnostics

Abstract
Biomedical datasets often contain a large sample imbalance and are subject to strict privacy constraints, which together hinder the development of accurate machine learning models. One potential solution is to generate synthetic images, as this can improve data availability while preserving patient privacy. However, it remains difficult to generate synthetic images of sufficient quality for training robust classifiers. In this work, we focus on the classification of single white blood cells, a key component in the diagnosis of hematological diseases such as acute myeloid leukemia (AML), a severe blood cancer. We demonstrate how synthetic images generated with a fine-tuned stable diffusion model using LoRA weights when guided by real few-shot samples of the target white blood cell classes, can enhance classifier performance for limited data. When training a ResNet classifier, accuracy increased from 27.3% to 78.4% (+51.1%) by adding 5000 synthetic images per class to a small and highly imbalanced real dataset. For a CLIP-based classifier, the accuracy improved from 61.8% to 76.8% (+15.0%). The synthetic images are highly similar to real images, and they can help overcome dataset limitations, enhancing model generalization. Our results establish synthetic images as a tool in biomedical research, improving machine learning models, and facilitating medical diagnosis and research.
1 Introduction
Accurate classification of white blood cells is essential to support clinical decisions and detect hematological disorders [18]. However, the current challenge in the classification of cytomorphological images is the lack of sufficient real data to effectively train models for rare classes [17]. It is widely acknowledged that the efficacy of a classifier is contingent on the quality and quantity of the training dataset [1]. The problem is that in many cases, obtaining an optimal dataset is not feasible, particularly when dealing with underrepresented conditions or infrequent cell types, where gathering sufficient data is expensive and time-consuming.
Furthermore, managing real patient medical data often requires strict privacy and authorization considerations, which can ultimately hinder access and collaboration across different institutions [10, 16]. An alternative is offered by synthetic data, which enables consistent data availability and facilitates collaborative research without compromising patient confidentiality [7]. Thus, high-quality synthetic images could serve as a reliable complement to real-world medical data [4].
In recent advancements of text-to-image generative models [11, 9, 2, 14, 13], diffusion models have emerged as a superior alternative in the domain of synthetic image generation. The capacity to produce high-quality images that accurately reflect the distinctive morphological characteristics of cellular structures renders them well-suited for medical applications [8, 12]. Furthermore, by effectively using diffusion models, it is possible to generate synthetic data that preserve the characteristic visual appearance and unique biological features of each cell type [8]. This opens the door to the possibility of obtaining unlimited data that reflect the morphological aspects of cells without privacy restrictions, potentially representing a major breakthrough in the biomedical field [15].
Real Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Synthetic Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Basophil Typical lymphocyte Monoblast Band neutrophil Promyelocyte bilobed
Figure 1: Comparison between real images (top row) and synthetic images (bottom row) for five different single white blood cell classes.
Despite recent advances, limitations still remain in cytomorphology, particularly regarding the generation of fine details for rare white blood cell classes with only few real samples, i.e., 4, 8, or 16. These limitations could be overcome with few-shot learning and generative models [5]. It presents strong potential, as simply using a few real samples could allow to train a model capable of generating a large amount of synthetic data. Instead of applying traditional image augmentations, i.e., rotations, brightness, or color-jittering, this approach is capable of generating new data that is not just a tuned copy of the originals, but novel instances that preserve biological plausibility, as illustrated in the comparison between real and synthetic single white blood cell images in Figure 1.
After the accurate synthetic image generation of single white blood cells, we train the classifier with solely high imbalanced real data, balanced synthetic data, and combination of real and synthetic data to show the effectiveness of the data generation step. To summarize, the contributions of our work are as follows:
  • We adapts Stable Diffusion model to generate highly accurate single white blood cell images for rare classes that further being used to train a better classifier for all single white blood classes.
  • Under the real few-shot guidance and customized text prompts with LoRA weights fine-tuning, the generated synthetic data align the best with the real images of rare classes.
  • We extensively evaluate the image generation and classification pipeline to validate the effectiveness of our pipeline (refer to sections 3.4 and 3.5).
2 Methodology
2.1 Single white blood cells image generation
The process begins with the generation of high-quality synthetic images that resemble real white blood cells in both visual and morphological aspects. The goal is to minimize domain shift and noise, ultimately improving classifier performance when trained on both real and synthetic data. In this work, the generative model, Stable Diffusion 2.1 [11], was used, a diffusion model trained on millions of images, capable of generating high-quality images from random noise. Once trained, it can receive a text description (prompt) and generate an image related to the description provided by the user. While Stable Diffusion 2.1 is capable of generating high-quality images, it frequently exhibits a lack of precision when attempting to produce a specific image, such as a white blood cell. The result is devoid of realism and the morphological characteristics necessary for use in diagnostic classification, as shown in Figure 2.
Figure 2: Synthetic eosinophil images generated with Stable Diffusion 2.1.
The most viable solution is to use the DataDream method [5], which consists of applying the Low-Rank Adaptation method (LoRA), a fine-tuning method to adapt a large pre-trained model.
The utilization of LoRA weights enables the implementation of minor yet effective alterations to a pre-trained model, such as Stable Diffusion 2.1 [3]. The efficacy of this approach is predicated on the incorporation of two low-rank matrices within the attention layers of the text encoder and the Diffusion U-Net. This modification facilitates the model’s ability to prioritize the most pertinent components of the prompt, thereby obviating the necessity for retraining of all layers.
The implementation of this technique results in a minor alteration to the image generation workflow. This approach necessitates the utilization of a prompt to generate synthetic images and the incorporation of LoRA weights that have been trained on real samples. The success of this methodology stems from the integration of text-to-image generation facilitated by Stable Diffusion 2.1 and image-to-image generation enabled by the trained LoRA weights, as illustrated in Figure 3.
Figure 3: We fine-tune LoRA weights with modified stable diffusion 2.1 model for the linear weights of the attention layers in both the text-encoder and the diffusion U-net encoder. We provide the few-shot real images to better learn fine-grained details of each classes to generate images closer to the target class. Abbreviations: Attn = Attention layer, FF = Feed-forward layer, Conv = Convolutional layer.
The initial step in the image generation process is to train the LoRA weights using real samples. Four different experiments were conducted per class, each using a different number of images: 1, 4, 8, and 16. It is very important that these images are as sharp as possible. To ensure this, a manual selection process was carried out to choose 16 images per class that clearly display the biological characteristics, since the dataset contains a wide range of image quality, as observed in Figure S1 of the supplementary material.
Subsequently, specific prompts must be written for each class. These prompts not only specify the type of blood cell but also include essential details that each cell type must exhibit. For each class, a trial-and-error study was conducted to select the best prompts. An example of a final prompt is the following:
”Photorealistic basophil under microscope, peripheral blood smear, large bilobed nucleus, dark blue-purple granules densely packed in cytoplasm, surrounded by red blood cells, medical cytology, high detail, clinical pathology, Wright-Giemsa stain, white background, soft lighting, 40x magnification, ultra-detailed, sharp focus, macro lens.”
2.2 Single white blood cells image classification
Subsequent to the generation of synthetic images, various experiments can be conducted using two different classifiers: ResNet-50 and CLIP.
Each classifier has been trained in different experiments: including the use of solely real images, solely synthetic images, and a combination of both real and synthetic images. This setup enables the evaluation of several aspects. By training only with real data, we can assess whether classifiers are capable of achieving good performance despite strong class imbalance, and establish a baseline for addressing the research question: demonstrate whether synthetic images, generated using few real samples, can improve the performance of classification models and support medical diagnosis. Training solely on synthetic images allows us to evaluate whether classifiers can recognize the morphological features present in synthetic samples. Finally, the mixed dataset scenario makes it possible to analyze whether the inclusion of synthetic data is beneficial in biomedical applications.
Two different model architectures were used in the classification experiments: ResNet-50 and CLIP, each representing a different paradigm in image understanding.
ResNet-50 is a convolutional neural network (CNN) with 50 layers and residual connections. It is widely used in computer vision tasks due to its ability to learn deep hierarchical features while avoiding vanishing gradients. In this project, ResNet-50 was adapted for multi-class classification by replacing its final fully connected layer with a layer matching the number of cell types. This architecture was used to learn morphological features from real and synthetic cell images in a purely supervised setting.
CLIP (Contrastive Language–Image Pretraining), on the other hand, is a dual-encoder model that jointly learns representations of images and text. In this work, the image encoder used was the ViT-B/32 variant, a Vision Transformer model, while the text encoder is based on a standard transformer architecture. Unlike traditional classifiers, CLIP aligns images and textual prompts in a shared embedding space. During the experiments, each test image was compared to the embeddings of the class-specific prompts, and classification was performed by selecting the class with the highest similarity. This architecture is particularly relevant for evaluating the consistency between the visual content of the image and the semantic information provided by the prompts.
By comparing both architectures under different data regimes; real only, synthetic only, and mixed, the study aims to assess the generalizability of synthetic data and its integration into traditional and prompt-based classifiers.
The primary loss function used in all experiments was cross-entropy loss, a standard choice for multi-class classification tasks. It quantifies the difference between the predicted class probabilities and the true labels, encouraging the model to assign high probability to the correct class.
When using only real or only synthetic images, the cross-entropy loss was computed over the entire training dataset in the standard way. In both cases, cross-entropy was used to optimize the models’ predictions, and AdamW was employed as the optimizer to adjust the weights through backpropagation. All experiments were trained for 100 epochs.
In mixed experiments, where both real and synthetic images were used, a combined loss strategy was implemented to control the contribution of each image type. A study was conducted to evaluate the best way to define the loss function, considering the different nature of the images. One of the strategies evaluated is outlined below:
$\text{Loss}_{\text{total}}=\lambda_{1}\cdot\left(\frac{n_{\text{real}}}{n}\cdot\text{Loss}_{\text{real}}\right)+(1-\lambda_{1})\cdot\left(\frac{n_{\text{synth}}}{n}\cdot\text{Loss}_{\text{synth}}\right)$
where $n=n_{\text{real}}+n_{\text{synth}}$ and $\lambda\in[0,1]$. This formulation allows adjusting the relative importance attributed to each type of image.
In the final strategy employed, the loss function was computed in a manner that treated both image types equally.
3 Experimental Results
3.1 Dataset details
The Munich AML Morphology Dataset [6] includes 18,365 expert-labeled single-cell images. The images are from peripheral blood smears of 100 patients without hematological malignancy and 100 patients diagnosed with acute myeloid leukemia. The patients were treated at Munich University Hospital between 2014 and 2017. The images were captured using the M8 digital microscope/scanner at 100x magnification with oil immersion, ensuring excellent resolution and distinct cell morphology. Pathological and non-pathological leukocytes were classified into a standard morphological scheme derived from clinical practice by trained experts. The dataset contains 15 blood cell classes with highly imbalanced sample counts, ranging from thousands of images in some classes to fewer than twenty in others, as illustrated in Figure 4. This imbalance provides an ideal scenario to evaluate whether our approach improves classifier performance, especially for minority classes.
Figure 4: Distribution of samples by class in Munich AML Morphology Dataset. This figure shows the number of images per class, sorted by decreasing order.
All the experiments underwent a essential step in the data, applying a 5-fold cross validation to ensure consistency and robustness in the results. For each of the folds, the data has been split as follows; 60% for the training set, 20% for the validation after each epoch, and 20% for the test set, to test the model with unseen data.
The training set comprises the majority of the images, the validation set was used to monitor overfitting and tune hyperparameters during training, while the test set was kept entirely separate and only used for the final evaluation of the models.
Due to the class imbalance inherent in the dataset, special care was taken to ensure the representation of minority classes in all splits. The training and validation datasets contained fewer images from minority classes compared to majority ones, which motivated the generation and incorporation of synthetic images to mitigate this imbalance.
All images were subjected to a standardized preprocessing pipeline prior to training. First, images were resized to a consistent resolution compatible with the input size requirements of the classifiers, specifically ResNet-50 and CLIP. Next, normalization was applied to scale pixel intensity values, enhancing the stability and convergence of the models during training.
To further improve model robustness, data augmentation techniques were applied on the training set, including random rotations, flips, and color jittering, simulating variations that occur in real microscopy images. No augmentation was applied to the validation and test sets to ensure unbiased evaluation.
3.2 Evaluation metrics
To evaluate the quality of the synthetic images generated during the experiments, the Fréchet Inception Distance (FID) was employed. FID measures the similarity between the distributions of real and synthetic images by comparing the statistics of features extracted from a pretrained Inception network [19]. Lower FID scores indicate higher fidelity and diversity of the generated images, reflecting their closeness to real data.
For assessing the performance of the classification models, several standard metrics were used to capture different aspects of their predictive ability. The overall accuracy was computed as the proportion of correctly classified samples among all predictions, providing a general measure of model correctness.
To better handle class imbalance and provide a more nuanced evaluation, the macro-averaged F1-score was calculated. This metric is the harmonic mean of precision and recall averaged equally across all classes, thus reflecting the balance between false positives and false negatives on a per-class basis.
Additionally, the Area Under the Receiver Operating Characteristic Curve (AUC-ROC) was computed in a one-vs-all scheme for each class. Unlike accuracy and macro F1-score, which rely on the final class prediction obtained via the argmax of the softmax outputs, the AUC-ROC metric works directly with the predicted class probabilities. This allows for evaluating the model’s ability to distinguish between classes across different decision thresholds, offering deeper insight into its discriminative power beyond the final decision.
Together, these metrics offer a comprehensive evaluation framework to measure both image generation quality and classifier effectiveness across the multi-class setting.
Table 1: Classification performance of ResNet-50 and CLIP models trained on 18,365 real images from 15 highly imbalanced blood cell classes. The dataset is split into 60% training, 20% validation, and 20% test sets. Results are averaged over 5 cross-validation folds.
Model (Split) Accuracy F1 macro AUC
ResNet-50 (Validation) 0.40 $\pm$0.07 0.08 $\pm$0.03 0.57 $\pm$0.07
ResNet-50 (Test) 0.27 $\pm$0.10 0.07 $\pm$0.01 0.56 $\pm$0.03
CLIP (Validation) 0.78 $\pm$0.01 0.27 $\pm$0.05 0.88 $\pm$0.02
CLIP (Test) 0.62 $\pm$0.03 0.23 $\pm$0.02 0.83 $\pm$0.02
3.3 Implementation details
The models were optimized using the AdamW optimizer, which combines adaptive moment estimation with decoupled weight decay (set to $1\mathrm{e}{-8}$) to promote better generalization. A cosine learning rate scheduler with warm-up was employed to adjust the learning rate dynamically across epochs. The learning rate was initially set to $1\mathrm{e}{-4}$, with a minimum value of $1\mathrm{e}{-8}$, and the warm-up phase lasted for 30 epochs.
Each model was trained for 100 epochs, using a batch size of 64 for training and 1 for evaluation.
3.4 Quantitative results comparison
3.4.1 Classification performance on real imbalanced data
Subsequently, an analysis of the metrics of the two classifiers utilized will be performed. The initial experiment was conducted employing real images solely as a point of departure for comparison and to determine our research question. Can synthetic images improve the performance of classification models and support medical diagnosis?
This outcome was expected under conditions of class imbalance. ResNet-50’s failure to consider minority classes has the effect of creating a false impression of high performance. As illustrated in Table 1, the accuracy achieved was 27%, which is four times higher than the performance of random guessing. However, examining the confusion matrix provided in the Supplementary Figure S2 reveals that it only correctly classifies samples exclusively from the Segmented Neutrophil, Typical Lymphocyte, and Monocyte classes, which are three of the most represented classes in our dataset.
In the scenario of imbalanced data, it is crucial to avoid reliance on accuracy metrics and to employ alternative metrics that are more reliable in such circumstances. One such metric is F1-Macro. This metric treats all classes equally, so it penalizes the model if it only classifies the majority classes correctly. The F1-Macro score of 0.075 indicates that the model is nearly random, rendering it incapable of differentiating between distinct cell classes and makes it obviously insufficient for clinical applications.
Concurrently, CLIP demonstrated superior performance. As illustrated in Table 1, the model achieved an accuracy of 62%, along with an F1-macro score of 0.23, which is approximately three times more accurate than the ResNet-50 model. The findings indicate that CLIP exhibits superior performance in scenarios characterized by significant class imbalance. This is due to the fact that, in contrast to receiving only images, CLIP receives a pair consisting of an image and the prompts for each class. This capability enables the system to meticulously analyze the morphological characteristics delineated in the prompts, thereby facilitating a more precise classification of the given data. Further insights into CLIP’s classification performance can be found in the confusion matrix shown in Supplementary Figure S3.
A notable observation is that both classifiers exhibit significant overfitting, as evidenced by the substantial discrepancy between the validation and test metrics. This could be because training for 100 epochs is excessive in this context. This causes the model to memorize the training data instead of generalizing.
Table 2: Classification performance of ResNet-50 and CLIP models trained on 45,000 synthetic images (3,000 per class) of 15 blood cell classes. The dataset is split into 60% training, 20% validation, and 20% test sets. Results are averaged over 5 cross-validation folds.
Model (Split) Accuracy F1 macro AUC
ResNet-50 (Validation) 0.93 $\pm$0.01 0.93 $\pm$0.01 0.99 $\pm$0.00
ResNet-50 (Test) 0.81 $\pm$0.01 0.82 $\pm$0.00 0.99 $\pm$0.00
CLIP (Validation) 0.87 $\pm$0.01 0.87 $\pm$0.01 0.99 $\pm$0.00
CLIP (Test) 0.77 $\pm$0.01 0.78 $\pm$0.01 0.98 $\pm$0.00
3.4.2 Classification performance on only synthetic data
Both classifiers were trained using exclusively 3,000 synthetic images for each class. This experiment is of significant value; it enables a comparison of the two classifiers with a perfectly balanced dataset. Conversely, it facilitates the evaluation of the generated synthetic images, as optimal classification outcomes signify that the classes possess distinctive and discernible morphological characteristics.
The results using synthetic images exclusively have been successful, demonstrating the best results observed in this project (see table 2). ResNet-50 achieved an accuracy and F1-macro score of 0.82, while the AUC was close to 0.99, indicating excellent class differentiation. However, the CLIP model performed slightly worse, suggesting that it did not interpret the cellular morphological details in the synthetic images as well as ResNet-50.
3.4.3 Classification performance on real and synthetic data
Ultimately, both classifiers were trained using exclusively 3,000 synthetic images for each class. This experiment is of significant value; it enables a comparison of the two classifiers with a perfectly balanced dataset. Conversely, it facilitates the evaluation of the generated synthetic images, as optimal classification outcomes signify that the classes possess distinctive and discernible morphological characteristics.
The results using synthetic images exclusively have been successful, demonstrating the best results observed in this project. ResNet-50 achieved an accuracy and F1-macro score of 0.82, while the AUC was close to 0.99, indicating excellent class differentiation. However, the CLIP model performed slightly worse, suggesting that it did not interpret the cellular morphological details in the synthetic images as well as ResNet-50. The detailed results of this experiment are shown in Table 2
Additional experiments conducted to supplement this study can be found in the Supplementary Material (see Figure S9 and others).
3.5 Ablation study
3.5.1 Varying the number of few-shot real images for synthetic image generation
Initially, an evaluation of the synthetic images generated using the diffusion model Stable Diffusion 2.1 with the implementation of LoRA weights in the Attention Layers has been conducted. The presence of morphological details in the images will be key for the optimal training of the classifiers. This is exemplified in Figure 5, which displays five rows, each containing four images. The initial row displays real Monoblast images from the Munich AML Morphology Dataset [6]. The subsequent four rows present synthetic images generated employing diverse LoRA weights. In the first instance, the LoRA was trained with a single real image; the second instance involved four images; the third instance involved eight images; and the final instance involved sixteen images. A clear progression can be observed as the number of real images used to train the LoRA increases, resulting in enhanced realism and similarity to real images.
In the first row, the cell is barely recognizable, with only basic features such as the cytoplasm and nucleus being identifiable. In the second row, a clear distinction between red blood cells and monoblasts is already apparent, as is typically observed in a peripheral blood smear. The final two rows exhibit a significantly higher degree of similarity to the authentic images, making them almost indistinguishable. Unique morphological characteristics of the monoblast can be observed, such as a rounded or slightly oval nucleus, a cytoplasm that appears more bluish than the nucleus, and the presence of fine cytoplasmic vacuoles.
Real images Refer to caption Refer to caption Refer to caption Refer to caption
1-shot Refer to caption Refer to caption Refer to caption Refer to caption
4-shot Refer to caption Refer to caption Refer to caption Refer to caption
8-shot Refer to caption Refer to caption Refer to caption Refer to caption
16-shot Refer to caption Refer to caption Refer to caption Refer to caption
Figure 5: Real and synthetic Monoblast images generated under different shot conditions.
Figure 6: Line plot comparing the classification accuracy (%) of ResNet-50 and CLIP models as the number of synthetic images per class increases. Each point shows the mean accuracy across 5 cross-validation folds, with error bars representing the standard deviation.
The supplementary material contains further examples of other cell types that follow the same structure (see, e.g., Figure S4 and Figure S5). It can be observed that most synthetic images appear darker than the real ones. This difference may be attributed to variations in illumination, staining, or other acquisition-related factors, which do not necessarily affect the morphological features relevant for classification.
The process of image generation has been successful for most of cell classes; however, certain classes show less satisfactory results, as illustrated in the comparison between real and synthetic images in Figure 1. A notable example of this issue is the Promyelocyte Bilobed. The generated images presented in Figure S6, are characterized by a lack of sharpness, which complicates their recognition. This may be attributed to two primary factors. The primary and most apparent aspect is the quality of the real images. Despite the implementation of a manual selection process to identify the best real images for training the LoRA weights, there are classes, such as Promyelocyte Bilobed, where the quality of these images is considerably lower in comparison to other classes. This phenomenon is explicable, as this particular cell type is uncommon. A promyelocyte with a bilobed nucleus is typically observed in patients diagnosed with acute myeloid leukemia.
Another potential factor is the Stable Diffusion 2.1 model may not have been trained on a specific type of image, such as a cellular anomaly. Instead, it is possible that the model has been trained on more common classes like neutrophils or lymphocytes.
3.5.2 Accuracy scaling by number of synthetic data
Thereafter, different experiments were conducted by progressively adding increments of 100 synthetic images per class to the existing real images, until reaching a total of 1,000 per class. The results, summarized in Table S1, present a notable enhancement in all metrics using ResNet-50. As more images are added, the classifier starts to correctly classify the minority classes, thereby aligning the accuracy and F1-macro score.
The results are very promising. To illustrate, in the experiment involving 700 synthetic images per class, both accuracy and F1-Macro exhibit a value of 0.60, while the Area Under the Curve (AUC) attains a value of 0.93.
In contrast to other metrics, AUC functions with probabilities rather than with a final decision. The model’s high AUC value, when compared to the other metrics, suggests that it is learning to distinguish the classes effectively in the representation space. However, its final prediction is not entirely accurate.
Given the continual enhancement of the metrics, the experiment series was expanded by incorporating an additional 2,000 and 3,000 synthetic images per class. A thorough examination of the confusion matrix from the previous experiment with ResNet-50, provided in Figure S7 in the supplementary material, reveals that the majority of the samples from all classes have been classified with a high degree of accuracy.
The same experiments were conducted using CLIP. As shown in Table S2 in the supplementary material, the results indicate that the classifier demonstrates a reduced capacity for enhancement when generating synthetic images. While the accuracy of ResNet-50 increases from 27% to 75%, the accuracy of CLIP rises from 61% to 73%. In both classifiers, overfitting was reduced as more images were used, indicating that the larger the dataset, the more epochs are required.
The primary factor contributing to the enhancement in the metrics is the gradual balancing of the dataset over time. When utilizing exclusively real images, the four majority classes constitute 95% of the dataset. However, when 3,000 synthetic images per class are incorporated, these four classes constitute slightly over 46% of the dataset. This modification balances the metrics of both classifiers, leading to a substantial enhancement in ResNet-50 and a minor improvement in CLIP. As observed in the confusion matrix provided in the supplementary material S8, CLIP can accurately classify the majority of samples, such as ResNet-50.
Given the persistent upward trend in metrics without any discernible indications of stabilization, two concluding experiments were conducted, utilizing 4,000 and 5,000 synthetic images per class. The Figure 6 illustrates the continued increase in metrics, suggesting a discernible advantage in relation to the number of images utilized.
4 Conclusion
This project has demonstrated that the generation of synthetic images using diffusion models such as Stable Diffusion 2.1, fine-tuned with LoRA weights, is a tool to be considered in the biomedical field. The quality of synthetic images is such that they exhibit cellular morphological features that closely resemble real ones, especially when high-quality real images are available.
From a classification perspective, it can be stated that the inclusion of synthetic data has led to a clear improvement in the metrics, thereby supporting the project’s research question: synthetic images can improve the performance of classification models and support medical diagnosis.
It has been demonstrated that, on imbalanced datasets, CLIP can understand morphological details of all classes better than ResNet-50 and classify minority classes. However, when synthetic data was included, ResNet-50 showed slightly better results. This is because the dataset is larger and the classes are more balanced. Ultimately, ResNet-50 proved to be more effective at interpreting and classifying synthetic images.
This approach has the potential to facilitate the unlimited and realistic generation of biomedical data, thereby addressing privacy-related legal concerns and logistical challenges. This methodology has the potential to expedite and improve clinical research, particularly for rare diseases.
In conclusion, this work establishes a substantial foundation for the consideration of synthetic images as a complementary tool in the biomedical field. In the long term, synthetic images could potentially replace real images in specific biomedical applications, helping to advance the fields of medical research and diagnosis.
Supplementary Material
The full supplementary material, including extended figures, tables, and additional experiments, is available at:
https://github.com/JanCarreras24/Final-Degree-Project/releases/tag/v1.0-TFG